Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Mark Dennehy 167 Applewood Hts Greystones Co. Wicklow Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Rules

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

NTSA PO Box 9 Blackrock Sorting Office Blackrock Co. Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

NSRA Lord Robert’s House Bisley Camp Brookwood Woking Surrey GU24 0NP United Kingdom

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

UCD Rifle Club PO Box 70 Students Forum UCD Belfield Dublin 4

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dave Cullen 37 Main Street Rathdrum Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Kealan Symes Ballard Cottage Ballinaclash Rathdrum Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dave Dunne 17 Templeview Place Clare Hall Dublin 17

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Gerry Campbell 4 Lohunda Park Clonsilla Dublin 15

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Walsh Carjon Lodge Cross Chapel Blessington Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

P.J. Harding 378 Ryevale Lawns Leixlip Co. Kildare

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Richard Slattery Kickham Barracks Clonmel Co. Tipperary

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Geoff Cooney Wilkinstown Navan Co. Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Tom Shalvey Dulane Kells Co. Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Norman Smith 12 Kingswood Park Belfast BT5 7EZ

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Robert Heaney 33 Brownlow Street Comber Co. Down

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Miss Agnes Graham 114 Towerview Cresent Bangor Co. Down BT19 6AX

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Joe Dible Cul-Cuile Enniskerry Road Co. Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Sandy Wilson 9 Abbey Court St. Fintan’s Park Monkstown Co. Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Martin Mace 29 Rossdowney Road Londonderry BT47 1PB

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Fitzmaurice 50 Cherry Grove Sallins Avenue Naas Co. Kildare

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Ivor Fletcher 127 Belfast Road Antrim Co. Antrim BT41 2AZ

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Mick Delaney 12 Hillside Park Rathfarnham Dublin 16

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Walls Conel Hill Balbriggan Co. Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Terry Wearen Duleek House Duleek Co. Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Patsy Tyrell 1 Shannon Hill Enniscorthy Co. Wexford

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Kieran Barry Coolmakee Fermoy Co. Cork

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John Walsh 18 Coolamber Road Rathcoole Co Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Robert McVeigh 86 Gilford Road Lurgan Co. Armagh BT66 7EB

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

M. Rowland Trinity Hall Rifle Club Secretary 46 Luker Avenue Henley-on-Thames RG9 2EU United Kingdom

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Glasgow University Rifle Club c/o Sport and Recreation Service Stevenson Building 77 Oakfield Avenue G12 8LT

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Trinity College (Hartford) Rifle Club c/o Trinity College 300 Summit Street Hartford CT 06106 USA

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Kundan Nepal Trinity College (Hartford) Rifle Club Captain Box 700894 Trinity College Hartford CT 06106 USA

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dr.A.M. Jory Hougue du Moulin Vale Guernsey GY3 5NG

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Tom McKnight 23 Meadow Drive Ballyclare Co.Antrim BT39 9XF

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

D. Hyde 86 Dunadrt Road Muckamore Antrim Co. Antrim BT41 4QJ

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Aberdeen University Rifle Club Butchart Recreation Centre University Road Aberdeen AB24 3UT

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Edinburgh University Rifle Club Pleasance Sports Centre The University of Edinburgh Old College South Bridge Edinburgh EH8 9YL

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Derek Wild Minorca Chapel Minorca Hill Laxey Isle of Man

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2001  30th May 2001

A Chara,

Entries are invited for the Trinity Top Ten 2001 Postal competition. This year is the 40^{th} anniversary of the founding of the modern DURC club and so for this year we are asking for a large number of entrants to help celebrate the anniversary. For this purpose, we have expanded the competition this year to contain three leagues. As well as the traditional 50m prone rifle league, this year will see the inclusion of a 25yard prone rifle league and a 10m Air Rifle league, as well as a seperate category in each league for Junior shooters (who must have been aged 18 or under on Janurary 1^{st} this year). A team event will not be held this year due to the workload increase anticipated from having two new leagues added to the competition.

Entry forms are attached to this letter. Completed entry forms and shot cards should be sent to the following address:

Mark Dennehy, DURC Secretary, Dublin University Rifle Club c/o DUCAC House 27 Trinity College Dublin Dublin 2 Ireland

An appropriate address label is included for the return of the entry forms : more will be sent out with the round nomination stickers. The entry fee is IR£5 per individual, and IR£1 for Juniors.

The completed entry form must reach the above address by no later than Friday June 22nd. For those competing from abroad, entry forms may be faxed to +353-1-677 2204 .

Nomination stickers can be sent out early to facilitate the inclusion of cards shot at calendar events: email Mark.Dennehy@cs.tcd.ie or fax the above number.

Common Rules

• A shooter may shoot for one club only in each league

• A shooter may enter as many leagues as he or she wishes

• A junior shooters score will not be counted for the seniors league.

• A round is prenominated by placing the appropriate sticker behind the bull of the first target.

• Only one sticker per card.

• The shooter should write his or her name clearly on the front of the target together with their entry number and the round number as obtained from the sticker.

• If the card is shot concurrently with another competition, the card must still be prenominated in the usual way and the card should be forwarded in the normal manner.

• ISSF regulations will be used as the basis for all technical resolutions. Ties however, will not be broken except for the final overall placing.

• A shooter who fails to return cards for two rounds will be disqualified from the competition. These rounds do not have to be consecutive.

• The coordinator will accept no responsibility for cards given to other D.U.R.C. members at shoots.

50m Prone Rifle League Rules

• This is an open competition, shot outdoors at 50m. Each round consists of 20 shots (one four-target card) fired during a twenty-minute detail.

25yd Prone Rifle League Rules

• This is an open competition, shot indoors at 25 yards. Each round consists of 20 shots (two ten-target cards) fired during two ten-minute details.

10m Air Rifle League Rules

• This is an open competition, shot indoors at 10m. Each round consists of 20 shots (four five-target cards) fired during four five-minute details.

Note that not complying with the above rules will lead to a penalty of one point (not gunscore) from that round.

Scoring

Each round is scored (points ex200). The entrants are then ranked in order on the basis of their score in that round. Points are awarded as follows. The individual with the highest gunscore will recieve the same number of points as there are entrants, the next placed shooted will receive one less and so on down. The cumulative sum of your points from the ten rounds will determine your overall placing. This method will be applied to all three leagues: the scores in each league will be independent of the other two leagues.

Due Dates

Note that these are the dates by which the cards must reach DURC, and not the date by which they must be shot. Cards may be shot and submitted ahead of time.

Entry Forms June 22 Round 1 July 13 Round 2 July 27 Round 3 August 10 Round 4 August 24 Round 5 September 7 Round 6 September 21 Round 7 October 5 Round 8 October 19 Round 8 November 2 Round 10 November 16

Results will be posted to entrants after each round is scored, and will also be placed on the D.U.R.C. Website at http://durc.tcd.ie . Any interim results available, and a report on the final results will be published in the NTSA newsletter, the Irish Marksman. It is hoped to announce the final results and present prizes at the DURC 10m Air Rifle Open on November 25^{th}.

Prizes

The highest placed ten shooters in each league will each receive a Trinity Top Ten patch. Medals will be awarded to the top three shooters in both senior and junior categories in each league, subject to entry levels. Good Luck to all !

Misé le meas,

Mark Dennehy, DURC Secretary